# FAST: An Efficient Scheduler for All-to-All GPU Communication

### v2

@@ 6 Related Work
Collective communication schedulers. Classic all-to-all algorithms such as SpreadOut [33] assume a single-tier network and balanced workloads. Modern GPU libraries like NCCL [3] and RCCL [5] exploit two-tier fabrics, but they are not designed to mitigate stragglers or incast under skew. Solver-based schedulers such as SCCL, TACCL, TE-CCL, and SyCCL [9, 44, 27, 10] use SMT or MILP to generate near-optimal schedules, but are fundamentally not scalable: scheduling can take minutes to hours, making them impractical for dynamic workloads. In contrast, FAST is the first scalable, on-the-fly scheduler for skewed, dynamic alltoallv in modern GPU clusters.
@@ 7 Conclusion
All-to-All communication is vital to modern distributed applications, especially MoE models. This paper presents FAST, the first polynomial-time, on-the-fly scheduler for skewed, dynamic All-to-All, which state-of-the-art schedulers struggle to handle. FAST exploits fast scale-up links to rebalance traffic and enforces balanced one-to-one transfers over scale-out, enabling efficient alltoallv communication. Evaluations on NVIDIA and AMD testbeds show that FAST outperforms state-of-the-art systems while reducing synthesis time by orders of magnitude.
This work does not raise any ethical issues.
@@ A.1 Performance Bound under Adversarial Workload
We first introduce the symbols and assumptions used for proofs. There are servers and GPUs and NICs within each server, making a total of GPUs participate in All-to-All. We denote the per-GPU bandwidth of scale-up and scale-out network as and . The scale-up network topology is a switch, while the other topology’s performance bound can be derived in a similar way. The total transfer size between two different servers and is denoted as while intra-server portion of All-to-All in server is denoted as . () and constitutes the complete All-to-All transfer workload. Note that does not represent anything and is thus set to be zero for the ease of writing proofs. We don’t prove the situation where the intra-node transfer is the majority of the All-to-All workload because there are more scale-out pairs than scale-up pairs in multi-node ML workload, making this scenario rare. So, we assume each server’s intra-node transfer size is no larger than the average of all the inter-node transfers, i.e., .
@@
In conclusion, under adversarial workloads, the worst-case performance gap of FAST relative to optimal is bounded by the scale-up to scale-out bandwidth ratio. With today’s hardware—for example, a 4-node cluster with 450 GBps scale-up on H100 [37] and 400 Gbps scale-out—this bound implies that FAST’s worst-case scenario completes within 2.12× of the theoretical optimum. In practice, this worst-case adversarial workload rarely happens and the performance is much closer to optimal as we show in the evaluation.

### v3

@@ 7 Related Work
Collective communication schedulers. Classic All-to-All algorithms such as SpreadOut [mpi_all2all] assume balanced workloads and single-tier networks. Modern GPU libraries (e.g., NCCL [nccl], RCCL [rccl]) exploit two-tier fabrics but do not handle skew, stragglers, or incast. Solver-based schedulers such as SCCL, TACCL, TE-CCL, and SyCCL [sccl, taccl, teccl, syccl] can produce near-optimal schedules but incur prohibitive overheads, making them unsuitable for dynamic workloads. In contrast, FAST provides a scalable, on-the-fly scheduler for skewed and dynamic alltoallv in modern GPU clusters.
MoE optimizations. Prior work on mixture-of-experts training [lina, tutel, gshard, switchtransformer, deepspeed_moe, mixtral, deepseek] improves efficiency through model design or communication–computation overlap, but typically treats alltoallv as a black box. FAST complements these efforts by directly optimizing alltoallv.
Leveraging NIC idleness. Recent work, FuseLink [fuselink], exploits under-utilized NIC bandwidth but operates below the collective layer, preserving the logical communication structure. Although FuseLink shifts traffic onto idle NICs, it does not reassign flow pairings and thus cannot address bottlenecks from skew or incast. FAST instead reshapes communication using overlay paths that stage transfers through intermediate GPUs, explicitly mitigating hotspots and contention.
Switch designs. Prior work applies Birkhoff’s decomposition to switch scheduling [birkhoff_switch1, birkhoff_switch2] and circuit-switched All-to-All communication [chronos]. FAST instead applies it at the GPU collective layer, running over commodity packet-switched networks without any switch changes.
@@ 8 Conclusion
All-to-All communication is critical to modern distributed systems. We present FAST, the first polynomial-time, on-the-fly scheduler for skewed and dynamic alltoallv. FAST absorbs skew using fast scale-up links and enforces balanced one-to-one transfers over scale-out, enabling efficient alltoallv communication. Across NVIDIA and AMD testbeds, FAST outperforms state-of-the-art systems while reducing synthesis time by orders of magnitude.
9 Acknowledgments
We thank our shepherd, Kai Chen, and the anonymous reviewers for their constructive feedback. We also thank Yonghao Zhuang, Zhihao Jia, Isabel Suizo, and Hugo Sadok for feedback on the drafts of this work. This work was supported by ONR Award N000142412059 and a Sloan Research Fellowship. Justine Sherry holds concurrent appointments as an Associate Professor at CMU and as an Amazon Scholar. This work was conducted at CMU and is not affiliated with Amazon. This work was also supported in part by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.
@@ A.1 Performance Bound under Adversarial Workload
We first introduce the symbols and assumptions used for proofs. There are servers and GPUs and NICs within each server, making a total of GPUs participate in All-to-All. We denote the per-GPU bandwidth of scale-up and scale-out network as and . The scale-up network topology is a switch, while the other topology’s performance bound can be derived in a similar way. The total transfer size between two different servers and is denoted as while intra-server portion of All-to-All in server is denoted as . () and constitutes the complete All-to-All transfer workload. Note that does not represent anything and is thus set to be zero for the ease of writing proofs. We don’t prove the situation where the intra-node transfer is the majority of the All-to-All workload because there are more scale-out pairs than scale-up pairs in multi-node All-to-All workload, making this scenario rare. So, we assume each server’s intra-node transfer size is no larger than the average of all the inter-node transfers, i.e., .
@@
In conclusion, under adversarial workloads, the worst-case performance gap of FAST relative to optimal is bounded by the scale-up to scale-out bandwidth ratio. With today’s hardware—for example, a 4-node cluster with 450 GBps scale-up on H100 [h100] and 400 Gbps scale-out—this bound implies that FAST’s worst-case scenario completes within 2.12× of the theoretical optimum. In practice, this worst-case adversarial workload rarely happens and the performance is much closer to optimal as we show in the evaluation.